The Collins Corporation purchased office equipment at the beginning of 2016 and capitalized a cost of $2,075,000. This cost included the following expenditures:

Purchase price $ 1,900,000
Freight charges 35,000
Installation charges 25,000
Annual maintenance charge 115,000
Total $ 2,075,000


The company estimated an ten-year useful life for the equipment. No residual value is anticipated. The double-declining-balance method was used to determine depreciation expense for 2016 and 2017.

In 2018, after the 2017 financial statements were issued, the company decided to switch to the straight-line depreciation method for this equipment. At that time, the company’s controller discovered that the original cost of the equipment incorrectly included one year of annual maintenance charges for the equipment.

Required:
1 & 2. Ignoring income taxes, prepare the appropriate correcting entry for the equipment capitalization error discovered in 2018 and any 2018 journal entry(s) related to the change in depreciation methods. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1

Depreciation Recorded: DDB = 2 x (1 ÷ 10) = 20%

2016 = 20% x 2,075,000 = 415,000

2017 = 20% x (2,075,000 - 415,000) = 332,000

Correct Depreciation:

2016 = 20% x 1,960,000 = 392,000

2017 = 20% x (1,960,000 - 392,000) = 313,600

Retained Earnings $ 73,600.00
Accumulated depreciation $ 41,400.00
Equipment $ 115,000.00

The change in depreciation method is treated as a change in estimate.

2018 Book Value = 1,960,000- 392,000 - 313,600

=1,254,400

Remaining Life = 8 Years

Straight Line Depreciation = 1,254,400/8 = 156,800

Depreciation expense $ 156,800.00
Accumulated depreciation $ 156,800.00